[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Jan 13, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Jan 13 12:44:52 UTC 2017


HSU Shift Log for Fri Jan 13 2017 07:44:43 GMT-0500 (EST)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: The western Pacific Himawari satellite imagery has  
been out since before the holidays. While it isn't mission critical, it is  
useful to have for global weather/climate monitoring.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170113/62f00038/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list